Let 𝐹 = 2π‘₯𝑧𝑖 βˆ’ π‘₯𝑗 + 𝑦2π‘˜. Evaluate 𝑉 𝐹𝑑𝑉 is the region bounded by the surfaces π‘₯ = 0, 𝑦 = 0, 𝑧 = π‘₯2 , 𝑧 = 4.


1
Expert's answer
2021-05-27T07:37:00-0400

Here,

"F= 2xzi-xj+y^2k"

The region V is covered

(a) by keeping x and y fixed and integrating from "z=x^2 \ to\ z=4" (base to top of column PQ),

(b) then by keeping x fixed and integrating from y=0 to y=6 (R to S in the slab)

(c) finally integrating from x=0 to x=2 (where"z=x^2" meets z=4)




Then the required integral is :


"\int\int\int_V\ Fdv = \int_{x=0}^2\int_{y=0}^6\int_{x^2}^4(2xzi-xj+y^2k)dzdydx"


"=i\int_0^2\int_0^6\int_{x^2}^4(2xz)dzdydx-j\int_0^2\int_0^6\int_{x^2}^4(x)dzdydx+k\int_0^2\int_0^6\int_{x^2}^4(y^2)dzdydx\\\ \\=128i-24j+384k"